Fortes Inc. has provided the following data concerning one of the products in its standard cost system. Materials price variances is calculated on material purchased and materials quantity is based on Material used in production. Variable manufacturing overhead is applied to products on the basis of direct labor-hours.

Inputs

Standard Quantity or Hours per Unit of Output  

Standard Price or Rate  

Direct materials

10.0

ounces

$

8.10

per ounce

Direct labor

0.5

hours

$

24.70

per hour

Variable manufacturing overhead

0.50

hours

$

6.40

per hour

The company has reported the following actual results for the product for April:

Actual output

7600   

units   

Raw materials purchased

50,200

ounces

Actual cost of raw materials purchased

$

386,210

Raw materials used in production

76,030

ounces

Actual direct labor-hours

3450

hours

Actual direct labor cost

$

88,880

Actual variable overhead cost

$

21,295

Required:

     Enter all numbers as positive and rounded to the nearest dollar   

Materials price variance

$

Materials quantity variance

$

Labor rate variance

$

Labor efficiency variance

$

Variable overhead rate variance

$

Variable overhead efficiency variance

$
Answer #1
1)
Direct Material
Quantity Rate Amount
Standard per unit 10 8.10 81.00
Standard 7600 units 76000 8.10 615600.00
Actual 7600 units 76030 7.70 585431.00
[Actual Rate of Raw Material Purchased: Actual Cost of raw material / Raw Material Purchased
= $386210/50200 Ounces
= $7.70 (Round off)
Direct Labor
Hours Rate Amount
Standard per unit 0.50 24.70 12.35
Standard 7600 units 3800 24.70 93860.00
Actual 7600 units 3450 25.76 88880.00
Variable Manufacturing Overhead
Hours Rate Amount
Standard per unit 0.5 6.40 3.20
Standard 7600 units 3800 6.40 24320.00
Actual 7600 units 3450 6.17 21295.00
Material Price Variance
It measures change in total cost on account of change in material cost per unit
= Actual quantity ( Actual Rate - Standard rate)
= 76030 (7.70 - 8.10)
= 30412 Favorable
Material Quantity Variance
It measures change in total cost on account of change in material quantity used
= Standard Rate ( Actual Quantity - Standard Quantity)
= 8.10 ( 76030 - 76000 )
= 243 Unfavorable
Labor Rate Variance
It measures change in total cost on account of change in Labor cost per hour
= Actual Hours ( Actual Rate - Standard rate)
= 3450 ( 25.76- 24.70)
= 3657 Unfavorable
Labor efficiency Variance
It measures change in total cost on account of change in Labor hours used
= Standard Rate ( Actual Hours - Standard Hours)
= 24.70 ( 3450 - 3800 )
= 8645 Favorable
similarly, rest of the variances will be calculated with the given formula:
Variable Overhead Rate Variance
= Actual Hours (Actual Rate - Standard Rate)
Variable Overhead Efficiency Variance.
= Standard Rate ( Actual Hours - Standard Hours)